# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v1

@@ III-B Simplification Process
The predicate filter eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity. Given a dataset of trajectories $\mathcal{D}$ and a predicate $\psi\in\Psi$, we define the robustness distribution as
@@
where
$\mathcal{D}_{1}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})\geq 0,\forall t^{\prime}\in t+I},$ (13)
@@
$\mathcal{D}_{3}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})<0,\forall t^{\prime}\in t+I}.$ (15)
@@
Let $\mathcal{T}_{p}=\{(\tau_{i},l_{i})|l_{i}=1\}_{i=1:N_{\mathrm{data}}}$ be the positive set of trajectories and $\mathcal{T}_{n}=T\setminus T_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we then calculate its robustness distributions for $\mathcal{T}_{p}$ and $\mathcal{T}_{n}$$d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, respectively—and their cosine similarity $S:=\frac{d(\mathcal{T}_{p},\psi)\cdot d(\mathcal{T}_{n},\psi)}{\lvert d(\mathcal{T}_{p},\psi)\rvert\lvert d(\mathcal{T}_{n},\psi)\rvert}$. As a result, predicates (and their negation) with a similarity score exceeding a user-defined threshold $S_{\mathrm{th}}$ are removed from $\Psi$ (and $\Psi_{\neg}$) prior to optimization.
[…]
$R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (16)
The intuition behind the max operation is to choose the larger sum of weights of a predicate and those of its negation. Next, we define disjunctive clause regularizer $R_{\mathrm{D}}$. Given a temporal clause with importance weights $\mathbf{w}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$, $R_{\mathrm{D}}$ is defined as,
$R_{\mathrm{D}}(\mathbf{w}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}w_{ik}w_{jk},$ (17)
where $w_{ij},w_{jk}\in\mathbf{w}$. The intuition behind the formulation is to differentiate the used predicates in a disjunctive clause in CNF from the predicates in the other disjunctive clauses within a temporal clauses.
Putting these regularizers together, the overall loss $\tilde{\mathcal{L}}$ is
~$\tilde{\mathcal{L}}(\mathcal{L},\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})=\mathcal{L}+\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})\\ +\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (18)
@@
where regularizer coefficients $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are small real values, and $\mathcal{L}$ is a desired loss function. In this work, we use the loss $\mathcal{L}$ proposed by [18],
$\mathcal{L}(\phi)=\sum_{j=1}^{N_{\mathrm{data}}}\exp(-\zeta l_{j}r(\phi,\tau_{j},t))$ (19)
@@
The weight pruning stage further simplifies the currently inferred wSTL specification by reducing the number of active predicates. This step proceeds in two phases. First, all weights with a value of zero are pruned, ensuring that these predicates are excluded from subsequent optimization steps. Second, among the remaining nonzero weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified parameter—are removed to eliminate the least contributing parts of the specification. After pruning, the remaining nonzero weights are normalized such that their sum equals one.
@@ III-C Trajectory Sampling
As described in Section II, the neural network is trained as a binary classifier to distinguish between positive and negative trajectories. Positive trajectories are obtained by executing the target policy in the environment of interest. Alternatively, they can be sampled from existing datasets, depending on the application setting. Negative trajectories are generated using a random-walk policy in this work, serving as contrasting samples for the classifier. However, the choice of policy used for generating negative trajectories is flexible and can be adapted based on the user’s objectives or the specific characteristics of the environment.
III-D Explanability Evaluation Metrics
Beyond classification performance, TLNet emphasizes the human-interpretability and alignment of inferred wSTL explanations with underlying policy behavior. While alignment is ultimately subjective and often requires qualitative assessment, we introduce three quantitative metrics—conciseness, consistency, and strictness—to characterize key aspects of explanation quality.
Let $\phi$ be an wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation. For example, our specifications expect $N=2$ temporal clauses. First, we define a conciseness metric as,
@@ III-D Explanability Evaluation Metrics
where $m_{n}$ is the number of predicates present in $n$-th temporal clause of $\phi$. Conciseness ranges from 0 to 1, with higher scores implying a shorter explanation with at least one predicate inside each expected temporal clause.
We now define a consistency metric. Let $\Phi=\{\phi_{k}\}_{k=1:K}$ be a set of $K$ explanations for the same policy (e.g., from different training seeds), and $\phi_{kn}$ be the $n$-th temporal clause ($n\in[1,N]$) expected to appear within $\phi_{k}$. We define a set of unique temporal clause structures $\Xi^{n}$ for a set of the $n$-th temporal clauses $\Set{\phi_{kn}}_{k=1:K}$. Each unique structure $\xi^{n}\in\Xi^{n}$ has $c_{\xi^{n}}$ temporal clauses satisfying $\xi^{n}$. Then, consistency of $\Phi$ is defined as
$\mathrm{Consistency}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\xi^{n}\in\Xi^{n}}\frac{c_{\xi^{n}}}{\lvert\Xi^{n}\rvert}&\text{if $\Xi^{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
Finally, we introduce a strictness metric. Let $\phi_{n}$ be the $n$-th temporal clause of an explanation $\phi$ and $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet). Denote the numbers of conjunctions and disjunctions in $\phi_{n}$ as $C_{n}$ and $D_{n}$, respectively. Then, strictness is defined as,
$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}$ is not empty},\\ 0&\text{otherwise.}\end{cases}$
@@ III-E Weight Distribution in wSTL
A key assumption in our neural network architecture is that weights can be distributed across Boolean and temporal clauses. We now prove that such distribution is theoretically sound. Recall that an arbitrary Boolean formula $\phi$ can be expressed in CNF as,
$\phi=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}(\neg)\psi_{j},$ (20)
where $n,m_{i}\in\mathbb{Z}_{>0}$ and $\psi_{ij}$ are predicates. The normalized weights of an arbitrary CNF wSTL formula can be syntactically distributed across Boolean operators as follows.
@@
$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{ij}^{\prime}(\neg)\psi_{ij},$ (21)
where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $w_{ij}^{\prime}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates. Furthermore, the weights $w_{ij}^{\prime}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}w_{ij}^{\prime}=1$.
@@ IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines: {outline} \1 A greedy pruning strategy proposed by [21], which independently removes individual network weights that do not affect classification accuracy post-optimization. \1 Top-$k$ pruning strategies (top-3 and top-5) used by [23, 20], which only retain the $k$ largest network weights. All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we aimed to explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively.
@@ IV-A Test Environments
We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied numbers of predicates.
@@ IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [16]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, roomba) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture aggressively targets the flag; fight pursues the blue agent; patrol defends a boundary region; and roomba switches between capturing and defending. In all red agent policies, the red agent performs random actions 25% of the time. We defined eight atomic predicates based on euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”) and binary defeat (“df”) status $\{1,-1\}$ of the red agent: $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.

### v2

@@ III-B Simplification Process
III-B1 Predicate Filter
As a pre-processing step, we implement a predicate filter that eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity. Given an arbitrary dataset of trajectories $\mathcal{D}$ and an atomic predicate $\psi$, we define the robustness distribution for $\psi$ as,
@@ III-B1 Predicate Filter
$\mathcal{D}_{1}$ :=$:=\set{\tau}{r(\psi,\tau,t)\geq 0\quad\forall t\in[0,H],\tau\in\mathcal{D}},$ (14)
@@
$\mathcal{D}_{3}$ :=$:=\set{\tau}{r(\psi,\tau,t)<0\quad\forall t\in[0,H],\tau\in\mathcal{D},}.$ (16)
That is, $\mathcal{D}_{1}$ is the set of trajectories where $\psi$ is always satisfied, $\mathcal{D}_{2}$ is the set where $\psi$ is satisfied at least once, and $\mathcal{D}_{3}$ is the set where $\psi$ is never satisfied.
@@
Now, let $\mathcal{T}_{p}=\{(\tau,l)|l=1,(\tau,l)\in\mathcal{T}\}$ be the positive set of trajectories in $\mathcal{T}$ and $\mathcal{T}_{n}=\mathcal{T}\setminus\mathcal{T}_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we then calculate $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions. We then remove predicates from the set $\Psi$ (and their negation from $\Psi_{\neg}$) which have a cosine similarity $S>S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
III-B2 Loss Regularization
The regularization step uses two regularizers to improve the conciseness of inferred explanations. The first regularizer, a temporal clause regularizer, $R_{\mathrm{T}}$, is defined as,
$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}s_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates.
The second regularizer, a disjunctive clause regularizer, $R_{\mathrm{D}}$, is defined as,
$R_{\mathrm{D}}(\mathbf{w}^{\mathcal{O}}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}\tilde{w}^{\mathcal{O}}_{ik}\tilde{w}^{\mathcal{O}}_{jk},$ (18)
where $\mathbf{w}^{\mathcal{O}}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$ defines the weights within one of our temporal operators and $\tilde{w}^{\mathcal{O}}_{ij},\tilde{w}^{\mathcal{O}}_{jk}\in\mathbf{w}^{\mathcal{O}}$. Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates.
Given $(\tau,l)\in\mathcal{T}$, our overall loss, $\tilde{\mathcal{L}}$, is then,
~$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
$+$ $\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (19)
@@ III-B2 Loss Regularization
where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters. We use $\lambda_{R_{\mathrm{T}}}=0.01$ and $\lambda_{R_{\mathrm{D}}}=0.1$ for this work. We use the classification loss from [21],
$\mathcal{L}(\tau,l,\mathbf{w})=\exp(-\zeta lr(\phi_{\mathbf{w}},\tau)),$ (20)
where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter ($\zeta=1$ for this work).
III-B3 Weight Pruning
@@
The weight pruning step further simplifies the currently inferred explanation by removing unnecessary weights. First, all weights with a value of zero are removed from the network. Then, among the remaining weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified hyperparameter—are removed to eliminate the least contributing parts of the specification. The remaining weights are then normalized. The pruning process terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the task-constraint structure of the inferred explanation.
@@ III-C Trajectory Sampling
We obtain positive trajectories in $\mathcal{T}$ by executing the target policy in the environment. These trajectories can also be sampled from existing datasets. We obtain negative trajectories in $\mathcal{T}$ by executing a random-walk policy in the environment.
III-D Explainability Evaluation Metrics
Beyond classification performance, TLNet also emphasizes the human-interpretability of inferred explanations. We quantify this emphasis by introducing three explainability metrics—conciseness, consistency, and strictness.
Let $\phi$ be a wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation. For example, in this work, we expect the inferred explanations to have one $\mathcal{F}$ clause and one $\mathcal{G}$ clause (thus $N=2$) because we assume there should be a task component and a constraint component. Now, let $\phi_{n}$ be the $n$-th expected temporal clause in $\phi$, where $\phi_{n}=\emptyset$ if that clause does not exist in $\phi$. Let the (unweighted) STL counterpart of $\phi_{n}$ be $\hat{\phi}_{n}$, where $n\in\{1,\dots N\}$. For example, in our work, if $\phi=\mathcal{G}(0.1\psi_{1}\wedge 0.9\psi_{2})$, then $\phi_{1}=\emptyset$, $\hat{\phi}_{1}=\emptyset$, $\phi_{2}=\mathcal{G}(0.1\psi_{1}\wedge 0.9\psi_{2})$, and $\hat{\phi}_{2}=\mathcal{G}(\psi_{1}\wedge\psi_{2})$.
III-D1 Conciseness
We define the conciseness of $\phi$ as,
@@ III-D1 Conciseness
where $m_{n}$ is the number of predicates present in $n$-th expected temporal clause $\phi_{n}$. Conciseness ranges from 0 to 1, with higher scores implying a shorter explanation with at least one predicate inside each expected temporal clause.
III-D2 Consistency
Let $\Phi=[\phi^{k}]_{k=1:K}$ be a sequence of $K$ explanations for the same policy (e.g., from different training seeds). Then, let $\Phi_{n}=[\phi_{n}^{k}]_{k=1:K}$ be the sequence of the $n$-th expected temporal clauses in $\Phi$, where $\phi_{n}^{k}$ is the $n$-th expected temporal clause for $\phi^{k}$. Let the STL counterpart of $\Phi_{n}$ be $\hat{\Phi}_{n}$. Then, let the set of unique elements of $\hat{\Phi}_{n}$ be $\Sigma_{n}=\{\hat{\phi}_{n}\in\hat{\Phi}_{n}\}$. We define the consistency of $\Phi$ as,
$\mathrm{Consistency}:=$
$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbf{1}_{\sigma_{n}}(\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise},\end{cases}$
where $\mathbf{1}_{\sigma_{n}}(\cdot)$ is the indicator function for $\sigma_{n}$. Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
III-D3 Strictness
Finally, let $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet). Denote the numbers of conjunctions and disjunctions in $\phi_{n}$ as $C_{n}$ and $D_{n}$, respectively. We define the strictness of $\phi$ as,
$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}\neq\emptyset$},\\ 0&\text{otherwise.}\end{cases}$
@@ III-E Weight Distribution in wSTL
A key assumption in TLNet is that the weights can be distributed across Boolean and temporal clauses. We now prove that such a distribution is theoretically sound. The normalized weights of a wSTL formula in CNF can be syntactically distributed across Boolean operators as follows.
@@
$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}\tilde{w}_{ij}(\neg)\psi_{ij},$ (21)
where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $\tilde{w}_{ij}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates. Furthermore, the weights $\tilde{w}_{ij}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}\tilde{w}_{ij}=1$.
@@ IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12): {outline} \1 NN-TLI [25]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization. We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Figure 2, since our empirical results showed more training stability with those functions. \1 STONE (top-3) [21]: this method retains the 3 largest network weights after training. \1 STONE (top-5) [21, 24]: this method retains the 5 largest network weights after training. All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively. Our code is available on GitHub after acceptance.
@@ IV-A Test Environments
We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied predicates.
@@ IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [19]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, mixture) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mixture switches between capturing and defending. Red agents perform a random action 25% of the time. We defined eight atomic predicates based on the defeated (“df”) status $\{1,-1\}$ of the red agent and euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”): $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.